3 px (1 point) Given It find the linearization of F at a = L(x) =
Answer #1

F(x) = 1 + 3 ha et at L(1) - Fla) +F(a) (1–0) FCO) = **tº dt = 70-4 F(x) = 0+3 . Al 1 etat] ;. 419 = F(3)+F(3)(x-3) #+ 3e\n-